OpenStudy (anonymous):

what is the limit of [[1/(3+x)]-(1/3)]/x as x approaches 0?

OpenStudy (lgbasallote):

\[\huge \lim _{x \rightarrow 0} \frac{ \frac 1{3+x} - \frac 13}x\] that's your question?

OpenStudy (anonymous):

yes. it looks a lot prettier now LOL

OpenStudy (anonymous):

multiply num and denum by 3(x+3)

OpenStudy (lgbasallote):

combine the fractions in the numerator first.. \[\huge \implies \lim_{x \rightarrow 0} \frac{\frac{3-(3+x)}{3(3+x)}}x\] do you get how that happened?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

good..so if you simplify that, you get... \[\huge \implies \lim_{x\rightarrow 0} \frac{\frac{-x}{3(3+x)}}x\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 0}\frac{ \frac{ 1 }{ 3+x } - \frac{ 1 }{ 3 }}{ x } = \lim_{x \rightarrow 0}\frac{ \frac{ 3+x - 3 }{ 3(3+x) } }{ x } = \lim_{x \rightarrow 0}\frac{ 1 }{ 3(3+x) }\]

OpenStudy (lgbasallote):

so simplify that further... \[\huge \implies \lim_{x\rightarrow 0} \quad \frac{-1}{3(3+x)}\] substitute x to 0 now...what do you get?

OpenStudy (anonymous):

maybe - is absent

OpenStudy (anonymous):

limit is -1/9

OpenStudy (lgbasallote):

right!!

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

omg THANKS!! I AM ALMOST DONE WITH TEST CORRECTIONS WOOO

OpenStudy (lgbasallote):

welcome